OpenStudy (anonymous):

If x= y-3 and y=z^2, what is x in terms of z?

OpenStudy (ajprincess):

Can u find y in terms of x @skittles223?

OpenStudy (anonymous):

no i dont know how not very good in math at all

OpenStudy (dmezzullo):

wat math class r u in?

OpenStudy (anonymous):

and if yall can give me step b step for my notes please

OpenStudy (anonymous):

im takin ged

OpenStudy (anonymous):

x=z^2 - 3 I think.

OpenStudy (dmezzullo):

geb?

OpenStudy (anonymous):

\[x=z ^{2}-3\]

OpenStudy (anonymous):

Not sure how to do step by step...

OpenStudy (anonymous):

Step-By-Step x=y-3 Substitute what why is into the equation Step completed

OpenStudy (anonymous):

y is*

OpenStudy (anonymous):

what is the other steps to find the answer

OpenStudy (anonymous):

That the only step there is :I

OpenStudy (anonymous):

wats y

OpenStudy (anonymous):

you told us what y is Y= Z^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so wats the answer

OpenStudy (anonymous):

@SchoolSlacker can you tell me answer i dont know how to figure it out

OpenStudy (anonymous):

I just said the answer in my first comment to this question

OpenStudy (anonymous):

x=z^2 - 3

OpenStudy (anonymous):

well they dont have that hold on i will show you the answers they have

OpenStudy (anonymous):

a. z^-6 b. z^6 c. z^-2/3 d. z^5 e. z^3/2

OpenStudy (anonymous):

well you should have put them on in the question alot more helpful for us

OpenStudy (anonymous):

answer is a its supposed to be x=y^-3